A 19 L container is filled with helium and the pressure is 143 atm and the temperature is 304 K. What volume can be filled in liters when the temperature is 28° C and the atmospheric pressure is 756 torr?


1
Expert's answer
2021-02-25T03:20:59-0500

n=pV/RT = 143 x 19/0.082 x 304 = 109 moles.

V= nRT/p= 109 x 0.082 x 301/1 = 2 690,338 L.